FEL Upgrade Project Weekly Brief
March 26-30, 2001

The FEL team successfully defended our progress on the design, analysis and construction activities for the 10 kW Upgrade project at this week's semiannual project review held on March 27.  As a useful
follow-up activity a 1-day workshop was held on FEL power scaling activities that took advantage of the review committee's presence.

According to the semiannual project review committee chaired by John Albertine, "there are no problems nor surprises" in our performance to date.  The committee noted its concern for missing budget elements that are presently not in the budget scope for FY02: (1) approx. $750k to finish the 3rd FEL Upgrade
cryomodule and commissioning funds. The committee also noted some concern for the design power loading levels (both peak and average) for the resonator optics.  A follow-up chart is being prepared for the committee that shows the upgrade design values are at or below power loading levels that have been already demonstrated in the 1 kW demo.

With the completion of the semiannual review, the project monthly report was completed for the month of February which includes a rebaselined cost-performance report that was shown and approved at the
review.  The rebaseline incorporates three engineering change request and the inclusion of the cost- performance plan for the Phase 2 of the project which extends from Feb.1, 2001 to Sept. 30, 2002.

We were pleased to have John Eric from the Air Force Research Lab attend the review.  John will be the contract monitor for the UV FEL portion of the Upgrade project in April which we anticipate will be
starting up sometime in April.

WBS 3 (Beam Physics):
We also started performance analysis on the IR Rev 1.1 design, which led immediately to work on optimization of the sextupole correction package. Further work was performed on beam loss instability analysis.

WBS 4 (Injector):
The design for the shield and internal cesiation scheme was completed, starting detailing. Ordered material to add weight into the FET system to allow for running at the higher fields expected with nitrogen implanted titanium. Ordering parts for a shield/mechanical movement test in vacuum.

Injector Gun HVPS - The spare Gun HVPS is at the vendor being upgraded.

WBS 6 (RF):
Zone 3 - No Progress.  Waiting for missing parts to arrive.

Zone 4 - The Mod Anode Current Monitoring system was repaired.  The cross-connect blocks were corrected and documented.  The Cavity Waveguide Window Temperature (CWWT) system was
tested and found to be working properly.  A door bracket is being made in the machine shop.  The
Circulator for #8 is still being repaired.  The documentation is still in drafting.

Injector RF - The Factory Acceptance Test for the first klystron has been scheduled for 6/20/01.  The construction of the Klystron Cart is in progress at the 10% level.  The Circulators are due 4/6/01.  Discussions are still ongoing with PSC concerning the upgrade of the Quarter's HVPS.

WBS 8 (I&C):
The rev. 3 design of the Optical Beam Position Monitor preamps is complete and is turned into EECAD. This version will be fitted into 2.5" aluminum box beam. The Molectron Q5 detector will be used, this is
5mm round split 4 ways. The magnification of the system will be reduced from 30 to 10. This will create a 1mm spot for the nominal 1 cm FEL beam, the system has a spatial dynamic range of greater than one full beam diameter in any direction. This version will be complete and ready for testing by the end of April.

James and Dan have wired the LSS box in Lab 2 and are nearly ready to power it up.  Additional work needs to be done on the LSS master prior to commissioning Lab 2.  Most of the parts are in for the user lab camera power supply chassis. Design is complete for the video patch panel.

The 3" diagnostics are beginning to arrive. This week the first shielded  beam viewers and BPMs came in. Procurements continue with the vacuum system components.  Discussions continue for a new design of the BPM electronics. Initial thoughts are centered on using the AD8313 logarithmic direct down-convertor (RF to DC), a sample/hold, and a low cost DSP. These would be located in a specifically designed compartment grouted into the new magnet stands. These devices would then be serially linked to the IOC, preferably with a commercial network like CAN (control-area-network).

Drawings in progress:
    Camera Power Supply wiring diagram was signed off.
    Picomotor control chassis is in final review, some has been signed.
    Charge/Dump current Monitor Chassis check print has been returned to EECAD for corrections and a
       wiring diagram has been added.
    New OBPM design is in for a drawing to be made.
    Fabrication Drawings for the Ion Pump Power Supply Front Panel were signed off and sent out for

Laser server has been up-graded.  Web work is on going.  Time has also been spent cleaning and re-stocking the I&C shop.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research continues to do small manipulations to their 3D model (in RADIA) of the GV dipole
    and has started the large Injector Dipole (GU).  The problem with the thin wedge magnet is that they
    can get good flatness without good integral gradient and vice versa.  They have not been able to get
    both properties within specification but continue to try different handles.  The start of GU analysis
    permits them to gain insight from a long magnet.
o  The drawing set details are being worked on for both magnets.  Any final adjustments to the design
    from the magnet analysis will have to be propagated back through the drawing sets.
Optical Chicane Dipole (DW)
o We finished the drawing set back-checking and are starting to work on the procurement specifications.
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing 180 degree dipole (GY) they added more length to the faces put
    a wedge groove into the path length corrector coil slot and widened the slot to the maximum available
    position as part of feed back from the magnetic model. They are now laying out the leads for the path
    length correctors.
o  The magnetic model will now be run with path length correctors on to test out their effect.
o The magnet steel specification remains in checking by Leigh Harwood.
QX (Trim Quad)
o  We continued to work on the procurement packages for this quadrupole and prepared to start pole tip
    chamfer studies.
o  We continued laying out the girders in 3D IDEAs format starting the placement of parts. The more
    complex regions more rapidly.
o  We started laying out the beam chamber in the arcs and will concentrate effort on them next week so
    the task parameters can be sent to AES.

WBS 10 (Wiggler):
We did a preliminary measurement of the dispersion section and are analyzing the data.
We are generating a task with AES to analyze the end coils of the Wiggler.

WBS 11 (Optics):
A portion of our week was spent giving presentations; to the Semi-Annual Review Team members, and to a panel convened to discuss our DoD-funded research contract.  As much in parallel as possible, we worked with designers on design of the gimbal mounts for the optical cavity assemblies, and on
the insertable mirror for the beam dump in the vault.  We received four 3" dia. uncoated, figured silicon substrates from the vendor one month earlier than promised.  Uncoated, they will be used in the
Mirror Test Stand with the Nd:YAG laser to provide power loading, and to test bonding techniques required to mate the mirror to the backplane assembly.  Once coated, they will be used in the upgrade optical cavity.

Preparations to ready User Lab 2 for cavity mirror tests continue.  The test pump down of the vacuum vessel late on 3/23 was successful.  The internal components to the mirror cassette were installed and aligned, as was the turning mirror above the cassette.  The I. & C. group continues installing cabling and
LSS hardware to support use of either this laser or the FEL in the lab.


No operations this week.